a. Explain the key role of a central bank (such as the Federal Reserve) in the...

  1.     a. Explain the key role of a central bank (such as the Federal Reserve) in the monetary system. (4 points).
  1. What happens to the money supply when a central bank (such as the Federal Reserve) buys bonds? Explain. (4 points).
  1. You run a bank. The current reserve ratio mandates holding reserves equal to 20% of deposits. If someone comes into your bank and deposits $10,000, by how much will the money supply in the economy increase? (4 points)
  1. You have equity (a capital share) in a bank. The bank has reserves, loans, and securities as well as deposits and capital. The bank suffers a sudden loss of value in one of its securities (assets). What will happen to the amount of capital the bank has and the leverage ratio? (4 points)
Answer #1

1) Answer: Monetary policies are the actions taken by central banks to control the monetary as well as financial status with the goal of attaining low inflation and sustainable growth in the economy. Broadly, there are two types of monetary policy namely - Expansionary and Contractionary. Expansionary monetary policy is used by FED when it wants to expands (increase) the supply of money to stimulate the economy. An expansionary monetary policy will stimulate investment and consumption spending, and also reduce interest rates thus leading to aggregate demand curve to shift right. Contractionary monetary policy is used by FED when wants to contracts (decrease) the supply of currency of the nation. This decreases the money supply, increases interest rates and reduces the aggregate demand. The tools that are used by the central bank under monetary policies include the reserve requirements, discount rate, and the open market operations.

2) Under the open market operations, the central bank buys or sells bonds to the public. In the scenario when the central bank wants to increase the money multiplier, the central bank buys bonds in the open market to increase the money supply which increases the money multiplier. When the central bank buys government bonds, thus putting money into the public, the demands for money supply will rise and the interest rates will be decline. Thus, it promotes investment spending and increases the quantity demanded of goods and services and GDP rises. In fact we can say that buying government bonds is a solution to prevent inflation

3) Reserve ratio = 0.2

Money multiplier = 1 / rr = 1 / 0.2 = 5

Thus money supply increases by $10,000 * 5 = 50,000